OpenStudy (anonymous):

4x+y=10 and 4x+2y=20 elimination method and show working out please

OpenStudy (anonymous):

y=10-4x y=(20-4x)/2=10-2x are you sure the question is right? the answer seems like zero =/

OpenStudy (anonymous):

4x + 2y = 20 (-)4x (-)y = (-)10 ----------------- y = 10 4x+y=10 =>4x=0 =>x=0